OpenStudy (anonymous):

Qstn related to Clocks.. How much does a watch lose per day, if its hands coincide every 64 minutes? plz help how to reach to solution..

OpenStudy (anonymous):

Do you have any idea on doing it?

OpenStudy (anonymous):

@Blaze : ya little bit i have done some sums..

OpenStudy (anonymous):

do you the answer?

OpenStudy (anonymous):

mean to say do you know the answer?check the answer sheet.

OpenStudy (anonymous):

Answer is 32(8/11) but i got 96 :(

OpenStudy (anonymous):

I also think it is 96 I got it too.

OpenStudy (anonymous):

@Blaze : but it's not the answer

OpenStudy (anonymous):

looks so and is tough

OpenStudy (anonymous):

@Ashleyisakitty

OpenStudy (anonymous):

what is the unit of the answer?

OpenStudy (anonymous):

1hour has 55min=>60/55(12/11)..to coinside 2 hands have to cover 60min=>12/11*60(720/11)..==.65.5/11.. bt clock coinside at 64min=>65.5/11-64=1.5/11=>16/11 min loss in 64min.. 16/11*1/64=1/44 min loss in 1min... in 24 hrs=>1/44*(24*60)=360/11=32.8/11 min the clock looses

OpenStudy (anonymous):

@shalinijagadeesan : nice one and wlcome to openstudy now put your own qstn and discuss here